OpenStudy (jessiegonzales):

medal!!! 2. -5y-9=-(y-1)

OpenStudy (jessiegonzales):

A. -1/2 B. -2 1/2 C. -2 D. -2/5

OpenStudy (anonymous):

Here we are trying to isolate y. What have you done so far?

OpenStudy (jessiegonzales):

SOLVE THE EQUATION

OpenStudy (jessiegonzales):

So far i just wrote down the problem i forgot where to start...

imqwerty (imqwerty):

-5y-9=-(y-1) -(5y+9)=-(y-1) 5y+9=y-1 4y=-10 y=-10/4 y=-5/2

OpenStudy (anonymous):

@imqwerty Don't give direct answers.. she didn't engage with you at all and you're just giving the blatant answer.

imqwerty (imqwerty):

sorry